The rectangular pool at the Peabody is 27 feet wide. To put border tile around the pool the contractor would need 134 feet of tile?

Answer by checkley79(3341)

P=2L+2W
W=27
134=2L+2*27
134=2L+54
2L=134-54
2L=80
L=80/2
L=40 ANS. FOR THE LENGTH.
PROOF:
134=2*40+2*27
134=80+54
134=134